OpenStudy (anonymous):

Solve 2x2 = 14x + 20

OpenStudy (anonymous):

4=14x+20 4-20 = 14x -16/14 = x -8/7 = x

OpenStudy (anonymous):

pls click the GOOD ANSWER button if satisfied witth the answer :)